Next Section | Table of Contents


Site 1173

Hole 1173B

At 0645 hr on 10 May we arrived on location, lowered the thrusters and hydrophones, and started making up five stands of drill collars. A positioning beacon was not immediately deployed because the scientists requested an offset of 50 m at 314° from the position of Hole 1173A, which was cored during Leg 190; the ship maintained station by using the GPS. The ship was over the desired position of Hole 1173B at 0810 hr on 10 May, and a seafloor positioning beacon (Datasonics SN2031, 14.0 kHz, 214 dB) was deployed at 0835 hr.

Reentry Cone and Casing Operations

At 1200 hr on 10 May, we began rigging up to run 20-in casing by welding a casing guide shoe to the bottom joint and then adding nine more joints of casing. This casing was then added to the short (pup) joint of 20-in casing attached to the bottom of the casing hanger that was already latched to the casing running tool; the total length of this assembly was 120.6 m. This assembly was then landed and latched into the reentry cone at 1845 hr; the running tool was unlatched so that the BHA could be run into the casing.

Next, we began assembling and running the drilling BHA into the casing. The BHA consisted of the 18.5-in bit, 22-in underreamer, bit sub with float, mud motor, and four stands of 8.25-in drill collars for a total length of 123.5 m. The casing running tool was attached to the top of the BHA; the BHA was lowered into and latched to the casing and reentry cone. The reentry cone, BHA, and casing were lowered through the moonpool at 2100 hr.

By 0800 hr on 11 May, we had lowered 4781.77 m of drill string and deployed the vibration-isolated television (VIT) camera in preparation for spudding Hole 1173B. A backup seafloor positioning beacon (Datasonics SN2029; 15.0 kHz; 214 dB) was deployed at 0635 hr. When the VIT camera reached the casing running tool at 1030 hr, we observed that the reentry cone was missing but that the running tool and casing hanger were still attached to the drill string. We began assembling the second reentry cone.

The VIT camera was recovered; we changed the VIT camera drill pipe guide sleeve with the larger casing guide sleeve so we could lower the camera over the casing to inspect it. At 1515 hr, the camera was lowered over the running tool, and we began inspecting the casing. The entire casing string, bit, and underreamer appeared to be intact. We continued to the seafloor with the drill string and camera to see if we could locate the lost reentry cone. We were unable to locate the reentry cone; the search was abandoned at 1930 hr; we began tripping out of the hole so we could rig up the second reentry cone.

On 12 May, we had raised the drill pipe, secured the casing in the moonpool, removed the casing running tool, racked the drill collars in the derrick, and laid down the mud motor, underreamer, and bit. At this time we had to perform the routine slipping and cutting of the drill line, which was completed at 1100 hr. The casing was taken apart at the first collar, and the second reentry cone was moved into position over the casing collar. The top piece of casing was reattached through the reentry cone; the casing was lowered and latched into the second reentry cone. This time we welded the casing hanger to the reentry cone to prevent it from detaching. We tripped the drilling BHA back into the casing and latched the running tool; the entire assembly with the second reentry cone was lowered through the moonpool at 1830 hr.

On 13 May at 0445 hr, we experienced a VIT camera problem but continued tripping drill pipe while the camera was changed out. The drill pipe was lowered to 4781.77 meters below rig floor (mbrf), the VIT camera was deployed again at 0600 hr, and we picked up the top drive and spaced out in preparation to spud. Hole 1173B was spudded at 0717 hr. We drilled in the 20-in casing with the underreamer and mud motor to 4926.1 mbrf (124.2 mbsf). With the 20-in casing shoe depth at 4922.5 mbrf (120.6 mbsf), we released the casing running tool at 1110 hr; we started tripping out of the hole with the drill pipe and VIT camera. The bit cleared the seafloor at 1225 hr and the rig floor at 2315 hr.

Logging While Drilling

As soon as the tools were laid down, we started assembling the LWD/MWD BHA at 0030 hr on 14 May. The Anadrill engineers loaded the radioactive logging sources at 0515 hr. We then picked up the top drive and tested MWD tools at 300 gpm (60 strokes per minute [spm]) and 800-psi standpipe pressure. By 0600 hr we had assembled and tested the 9.875-in bit, LWD bit sub with float, LWD RAB tool, MWD tool, LWD ISONIC tool, LWD azimuthal neutron and density tool, and the LWD crossover sub. We added five drill collars, a tapered drill collar, and six joints of 5.5-in transition drill pipe, for a total length of 139.88 m. We removed the top drive and calibrated the Anadrill drawworks encoder. At 0700 hr on 14 May, we started tripping the drill string down to the seafloor. Unfortunately, we had to stop at 0930 hr (at 1406.81 mbrf) because of an approaching tropical storm (Cimaron). We tripped out of the hole and secured for transit by 1615 hr on 14 May to avoid the approaching storm.

We continued the transit to evade Tropical Storm Cimaron until 0030 hr on 15 May, when the captain felt it was safe to return to the site. The ship arrived back on site at 1100 hr on 15 May. The total transit was 185 nmi and lasted just over 18 hr.

Once we were back on location, the LWD/MWD BHA was assembled, the radioactive sources were loaded, and the tools were tested. We began lowering the LWD/MWD tools to the seafloor at 1330 hr (15 May). When the drill string reached 3876.06 mbrf, the VIT camera was deployed. However, the camera signal was lost when it reached 102 mbrf. The camera was retrieved, and we found that the camera cable had been damaged in the upper guide horn. Approximately 110 m of cable was cut off, and the cable was reterminated. While this was being done, the drill string was tripped to 4769.57 mbrf. The repaired VIT camera was deployed at 0330 hr on 16 May. While the camera was being lowered to the seafloor, the drill crew assembled extra drill collar stands in preparation for the next site.

The VIT camera reached the top stabilizer/sensor pad on the LWD tool at 0830 hr but could not be lowered past it. The installed VIT guide sleeve (the same one used for the drilling tools) had an internal diameter of 9.5 in, and the external diameter of the stabilizer pad was 9.875 in. The camera was retrieved, the guide sleeve changed to the casing guide sleeve (internal diameter of 27.5 in), and the camera was redeployed at 0830 hr (16 May). The reentry cone came into view at 1002 hr and was reentered in 1.5 hr. The drill string was lowered into the casing while the VIT was retrieved.

The drill string was spaced out in preparation for drilling and the Anadrill LWD/MWD tools were tested. Drilling/logging operations began at 1500 hr on 16 May at 4922.49 mbrf (120.49 mbsf). The drillers attempted to maintain a 50 m/hr rate of penetration (ROP), as the scientists had requested. The effective ROP was lower than this due to the time required to make pipe connections. The average surface drilling parameters were a WOB of 4–7 kilopounds (klb) at 60 rpm, a torque of 150–175 A, and a flow rate of 327 gpm (67 spm) with a standpipe pressure of 1250 psi.

LWD/MWD operations continued until 1545 hr (17 May) when Hole 1173B reached a total depth of 5539.00 mbrf (737.10 mbsf). The top drive was disconnected from the drill string and racked back in preparation to pull out of the hole. Due to deteriorating hole conditions, however, the top drive had to be reattached to provide rotation and allow circulation. The maximum overpull was 30 klb. The top drive was removed at 5346.6 mbrf. During the trip out of the hole, the LWD tools were recording the entire interval and the speed was controlled at 70 m/hr from 5221.90 to 5171.90 mbrf (420.00–370.00 mbsf) to evaluate LWD ISONIC velocity data quality. The drill bit cleared the seafloor at 2135 hr on 17 May.

Heave Compensation Experiments

Two experiments were conducted during the drilling of the hole to assess the efficacy of the shipboard active heave compensation (AHC) system. We used the MWD tool to measure downhole drilling parameters, including downhole WOB, torque, and bit bounce. These measurements are made using paired strain gauges near the base of the MWD collar, are transmitted to the surface for recording, and are not in downhole memory. Strong MWD pressure signals were recovered at the surface using MWD pulse rates of 3 bits per second (bps) at up to 5500 mbrf depth in Hole 1173B. Surface data from the rig instrumentation system (RIS), including surface rotations per minute, torque, WOB, ship heave, pitch, and roll, were recorded synchronously. Sea states were recorded up to 2 m and heave to 1.5–2 m during drilling at Site 1173.

For the first experiment six intervals were drilled with the AHC off (4955.7–4974.66, 5080–5099.29, 5224.15–5243.31, 5387.65–5406.87, 5464.62–5493.94, and 5527–5531.98 mbrf). During these intervals, drilling proceeded using the passive heave compensation system that has been used for most ODP drilling prior to Leg 189. The second experiment consisted of drilling four intervals with the AHC preloaded, which allows the driller to set the surface weight prior to putting the bit on the bottom of the hole (5329–5358, 5406–5435, 5435–5464, and 5522–5527 mbrf). Whether or not this practice is effective in controlling penetration rates is debatable.

Real-time shipboard observations suggest that the AHC damps high-frequency variations in uphole rotations per minute and torque, as well as the high-frequency variation in downhole WOB. The comparison of downhole MWD parameters with the surface information will be analyzed postcruise to evaluate the shipboard heave compensation system and drilling practices.

Hole 1173C Logging While Drilling

Hole 1173C was designed to obtain LWD/MWD logs from the uppermost section that was cased off in Hole 1173B. The ship was offset 50 m to the northwest of Hole 1173B, and Hole 1173C was spudded at 2320 hr on 17 May. The LWD/MWD drilling/logging operations were conducted from the seafloor (4801.90 mbrf) to a total depth of 4976.90 mbrf (175.00 mbsf), which was reached at 0500 hr on 18 May. The ROP was controlled at 60 m/hr and the effective ROP (including the time required to make pipe connections, etc.) was 30.9 m/hr. The average surface drilling parameters were a WOB of 4 klb at 60 rpm, a torque of 150 A, and a flow rate of 327 gpm (67 spm) with a standpipe pressure of 1250 psi.

The drill string cleared the seafloor at 0650 hr on 18 May; the LWD/MWD BHA reached the rig floor at 1430 hr. The LWD/MWD tools were disassembled, and the drill bit cleared the rotary table at 1655 hr on 18 May. One of the two seafloor positioning beacons was retrieved, and the other was turned off and left on the seafloor for use later in the leg when we would conduct ACORK operations. The drilling equipment was then secured for the transit to the next site.

Next Section | Table of Contents